3. A SpaceX StarshipTM single-stage rocket is launched vertically from the moons surface. It burns k kg of fuel per second, s(b) Integrate again to show that the altitude versus time is given by (e) = vast = 391 _ (20 – keybex in (m ort)
Answer #1

(a) By This figure a f = Fy - mg - V fama - fx = Ver dm = K Ven =) So, Equaticm One become I ma = kVex - mg a= k Vex -g Im=mo

yle) a) (b) Vit) = so, form] = fueldt ye): ver In [mo:] - gre) dt 407 = -filen en (mat) dt - faldt yes = = kan en - (243. Letlo-kt mo Sre) mouse please em (main) (n=1293-C0-13] | muon corak) (no titl - (1-44) + ]*g maken [. -*t) en (me ) - 1 + kit +